# Little By Little: Continual Learning via Incremental Mixture of Rank-1 Associative Memory Experts

### v3

@@ 3.3.2 Self-Activation for MoRAM Routing
Relying on the key-value form of the rank-1 adapter, we propose a self-activation mechanism for routing the memory experts. We derive the mixing weights directly from the intrinsic activation of each atomic memory, leveraging the key vectors $\mathbf{A}_{i,:}$ for memory retrieval. Unlike standard MoE approaches that rely on additional routing networks, our self-activation routing performs routing as a content-addressable retrieval, reducing the possibility of forgetting caused by an additional router.
Self-activated relevance scoring. By eliminating the external router, MoRAM avoids the optimization instability and forgetting associated with auxiliary networks. Instead, we derive mixing weights directly from the intrinsic alignment between the input and the memory keys. Formally, for an input token $\mathbf{x}\in\mathbb{R}^{d_{\text{in}}}$ and the set of accumulated memory atoms up to task $t$, we compute the raw relevance score ${\textnormal{s}}_{i}$ for the $i$-th atom as:
@@ 3.3.4 Continual Learning with MoRAM
Incremental memory expansion. Unlike methods that add fixed-size adapter blocks [96], MoRAM expands memory “little by little.” For each new task, we introduce a set of $r$ new atomic rank-1 pairs (Keys $\mathbf{A}$ and Values $\mathbf{B}$), freeze all prior atoms, and let the self-activation mechanism jointly route across the union of all old and new memories. This seamlessly integrates new knowledge while preserving the reusability of prior tasks. To enhance efficiency and enable sub-linear expansion rate, MoRAM supports low-utility atoms to be pruned after training with minimal performance degradation (more in Appendix A.10).
@@ A.10 Computation Cost
Analysis of post-pruning. While this approach entails linear storage growth, our atomic structure offers a distinct advantage over conventional LoRA. Unlike standard matrices that require complex post-hoc decomposition (e.g., SVD) to compress, our rank-1 experts are independent and can be individually assessed. Consequently, the model inherently supports post-pruning, retaining only atoms that exceed a cumulative activation threshold, allowing for storage reduction with minimal performance degradation if required.

### v4

@@ 3.3.2 Self-Activation for MoRAM Routing
Relying on the key-value form of the rank-1 adapter, we propose a self-activation mechanism for routing the memory experts. We derive the mixing weights directly from the intrinsic activation of each atomic memory, leveraging the $\mathbf{A}_{i,:}$ of each rank-1 adapter component, instead of introducing a separate router for the pairs of $\mathbf{A}_{i,:}$ and $\mathbf{B}_{:,i}$ in a standard MoE-LoRA modeling. It is naturally the key vectors of the memory modeling for memory retrieval. Unlike standard MoE approaches that rely on an additional routing module, our self-activation routing performs routing as a content-addressable retrieval, reducing the possibility of forgetting caused by an additional router.
Self-activated relevance scoring. By eliminating the external router, MoRAM avoids the additional optimization instability and forgetting associated with auxiliary network parameters (i.e., the router). Instead, we derive mixing weights directly from the intrinsic alignment between the input and the memory keys. Formally, for an input token $\mathbf{x}\in\mathbb{R}^{d_{\text{in}}}$ and the set of accumulated memory atoms up to task $t$, we compute the raw relevance score ${\textnormal{s}}_{i}$ for the $i$-th atom as:
@@ 3.3.4 Continual Learning with MoRAM
Incremental memory expansion. Unlike methods that add fixed-size adapter blocks [96], MoRAM expands memory “little by little.” For each new task, we introduce a set of $r$ new atomic rank-1 pairs (Keys $\mathbf{A}$ and Values $\mathbf{B}$), freeze all prior atoms, and let the self-activation mechanism jointly route across the union of all old and new memories. This seamlessly integrates new knowledge while preserving the reusability of prior tasks. To simply enhance efficiency and enable sub-linear expansion rate, MoRAM supports low-utility atoms to be pruned after training with minimal performance degradation (more in Appendix A.10), which can be further enhanced with more advanced dynamic expansion strategies.
@@ A.10 Computation Cost
Analysis of post-pruning. In the default implementation setup, MoRAM increments memory units with a constant number for each task, which entails a linear growth rate for the memory size (i.e., number of experts or model size). To enhance the efficiency and enable a sub-linear expansion rate, a more advanced dynamic expansion strategy [78] can also be applied jointly with MoRAM. In simple terms, after incrementing $r$ rank-1 atomic experts and training them, low-utility components can be identified and pruned, achieving a sub-linear expansion rate.
@@
The atomic design of MoRAM enables fine-grained modeling and flexible capacity expansion compared to conventional MoE–LoRA. In standard MoE–LoRA, experts are coarse-grained (i.e., LoRA adapters with higher rank), so adding or removing capacity is restricted to operate at a relatively large unit, or it requires post-hoc decomposition (e.g., SVD) or compression.